OpenStudy (anonymous):

-y - 2 = x + 1 is a decreasing function

OpenStudy (raffle_snaffle):

make it a function of y respect to x

OpenStudy (abb0t):

Yes. If you put it in the form: \(\sf \color{red}{y=mx+b}\) You will see

OpenStudy (anonymous):

ok so it is increasing y=-2x+1

OpenStudy (raffle_snaffle):

what makes you say it's increasing?

OpenStudy (anonymous):

sorry decreasing

OpenStudy (abb0t):

Yes.

OpenStudy (raffle_snaffle):

okay right!

OpenStudy (anonymous):

god help us it is not your fault, but \[-y - 2 = x + 1\] is not a function at all it is an equation

OpenStudy (abb0t):

It's a linear function, no?

OpenStudy (anonymous):

I have it in my homework it says if it is increasing or decreasing

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no it is not a linear "function" it is an equation for a line thus beginning the long confusion between the concept of "function" and "equation" it is not your fault just bad math teachers

OpenStudy (anonymous):

aaaa

OpenStudy (anonymous):

anyway what they want you to do is solve for \(y\) \[-y - 2 = x + 1\] \[-y=x+3\] \[y=-x-3\]

OpenStudy (anonymous):

but is it increasing or decreasing

OpenStudy (anonymous):

and then say "since the slope is negative, it is decreasing

OpenStudy (anonymous):

\[y=-x-3\\y=mx+b\] \[m=-1\] so it is decreasing

OpenStudy (anonymous):

ok thank you. was not sure

OpenStudy (anonymous):

yw sorry you have such a crappy book

OpenStudy (anonymous):

y = -5x + 4 so this one is increasing right

OpenStudy (anonymous):

i meant decreasing sorry didnt see the -

OpenStudy (anonymous):

yeah, decreasing

OpenStudy (anonymous):

k thanks